OpenStudy (anonymous):

7. Jeremy is building a large deck for a community center. The deck is shaped as a rectangle. The width of the deck is 29 feet. The perimeter of the deck is to be at least 134 feet. a. Write an inequality that represents all possible values for the length of the deck. b. Find all possible values for the length of the deck.

OpenStudy (hunus):

The perimeter of a rectangle is 2 times its width plus 2 times its length \[\large P = 2w+2l \] So if the perimeter has to be 'at least' 134 feet what do you think the inequality will look like?

OpenStudy (hunus):

If the perimeter were going to be equal to 134 what would the equation be?

OpenStudy (anonymous):

so... 134 = 2 x 33.5 + 2 x 33.5

OpenStudy (hunus):

We know the width is 29 feet

OpenStudy (anonymous):

134 = 2x29 + 2 x 37.5?

OpenStudy (hunus):

Let's just leave it as \[134 = 58 + 2l\] Now that equal sign should be <, <=, >,= or >= The perimeter should be at least 134 so which should it be?

OpenStudy (anonymous):

> the perimeter would be dubble l+W

OpenStudy (hunus):

Here, 58 + 2L represents the actual perimeter of the fence and 134 represents what the perimeter should be at the least

OpenStudy (hunus):

So the options would be 58 + 2L > 134 58 + 2L >or= 134 58 + 2L < 134 58 + 2L <or= 134 58 + 2L = 134 Based off of them saying that it should at the least be 134 which inequality do you think it is?

OpenStudy (hunus):

Can the perimeter of the fence be less than 134?

OpenStudy (anonymous):

58 + 2L >or= 134

OpenStudy (hunus):

Yep :D

OpenStudy (anonymous):

sorry son was srying

OpenStudy (hunus):

That's fine

OpenStudy (hunus):

Now isolate the L in 58 + 2L >or= 134

OpenStudy (anonymous):

134-58 = 76 2l<76 l<76/2 l<38

OpenStudy (anonymous):

not sure what it means by finding all possible values.

OpenStudy (anonymous):

All possible values for l with the W of 29 are all the values that are greater than 38.

OpenStudy (anonymous):

figured it out thanks hunus

OpenStudy (hunus):

Yup :)